OpenStudy (sarahc):

Each dimension of a rectangular prism is increased by 80%. What percent of the original volume is the volume of the new prism? 51.2 percent 324 percent 583.2 percent 64 percent

OpenStudy (sarahc):

I got 583...am I correct?

OpenStudy (anonymous):

Rectangular Prism = Length*Base*Height = (9/5)^3 LBH = 729/125 = 5.832.. Yes you are right..

OpenStudy (lgbasallote):

let's assign some values to make it easier volume = length x width x height let's say length = 2 width = 3 height = 4 if i increase by 80% length = 3.6 width = 5.4 height = 7.2 new volume = 139.968 this is 583.2% of the original volume so you're right

OpenStudy (sarahc):

Yay. just checking. thank you

OpenStudy (sarahc):

one more question please? Part 1: Create and provide the dimensions for two similar figures of your choosing. Part 2: What is the similarity ratio of these figures along with the ratio of their surface area and volume? Part 3: Show your work, either using the actual volumes or using the formula, that the volume ratio is true.

OpenStudy (sarahc):

igbasallote!